The following facts apply to a convertible bond making semiannual payments:

Conversion price $ 46 /share

Coupon rate 7.2 %

Par value $ 1,000

Yield on nonconvertible debentures of same quality 8 %

Maturity 30 years

Market price of stock $ 45 /share

What is the minimum price at which the convertible should sell?
Answer #1

Price of straight bond=1000*7.2%/8%*(1-1/1.08^30)+1000/1.08^30
=909.9377333

Price of converted shares=1000/46*45=978.2608696

Minimum price=978.2608696